OpenStudy (anonymous):

If you toss a coin 4 times, what's the probability that you'll get 3 heads?

OpenStudy (xapproachesinfinity):

think of all the possible arrangement that can occur when we toss a coin four time HHHH HHHT HHTT HTTT TTTT THHH and so on

OpenStudy (xapproachesinfinity):

count all the possibilities and count how many of them are 3 heads

OpenStudy (xapproachesinfinity):

there is an easier way of doing this but it is bit difficult to grasp why we are using the concept so go with the basic idea of counting

OpenStudy (anonymous):

Sir find probability just

OpenStudy (anonymous):

I am basically not sure if it is correct!!!!!

OpenStudy (anonymous):

1/2 is the probability

OpenStudy (ybarrap):

The probability of getting 3 heads and one tail in 4 tries is, by the Binomial Theorem $$ {4\choose3}\left ({1\over2}\right )^3\left ({1\over2}\right)^1 $$ See - http://en.wikipedia.org/wiki/Binomial_distribution#Probability_mass_function

OpenStudy (mathstudent55):

There are 2 * 2 * 2 * 2 = 16 total number of possible outcomes. Of those, only the following 4 outcomes have exactly 3 heads. HHHT HHTH HTHH THHH probability = (number of desired outcomes)/(total number of outcomes)